# Time-Aware Auto White Balance in Mobile Photography

### v1

@@ 5 Conclusion and limitations
While our method represents a promising solution for mobile camera ISPs, its dependency on contextual metadata limits its optimal accuracy to devices that provide geolocation data, which may not available on most DSLR cameras. Additionally, while the contextual metadata is device-independent (i.e., the differences across devices are expected to be minimal), our method relies on additional capture information (i.e., ISO, shutter speed, noise, and SNR stats) and image colors, making it inherently camera-dependent in design. This dependency prevents our trained model from generalizing to new devices without fine-tuning or re-training. However, this issue can be mitigated through color calibration (e.g., [50]), which can be extended to calibrate capture information as well—by performing a pre-processing mapping from the new camera space (for both color and capture information) to the camera space used during training.
@@ Appendix B Additional details and results
Additional details on contextual information: In the main paper, we presented our method, which relies on the “probability” of an image being captured during one of the key solar events (e.g., sunrise, noon, sunset) along with additional capture metadata and color information represented as histograms.
@@
Additional ablation studies: In the main paper, we presented a set of ablation studies to analyze the impact of different input features on our method. Here, we provide additional ablation studies on the validation set with masks applied, as shown in Table 5. By default, our results in Table 5 use the histogram feature, $\mathbf{H}$, and the time-capture feature, $\mathbf{c}$, with noise stats, $\mathbf{n}$.
@@
Analysis on outdoor vs. indoor scenes: For outdoor scenes, time-of-day information provides strong cues about the likely range of illuminants. However, it is not as informative for indoor scenes, which are typically illuminated by artificial lights. Therefore, for indoor scenes, image colors and additional capture information are necessary for accurate illuminant estimation.
@@
Additional quantitative results: In the main paper, we reported results on our testing set without masking out regions illuminated by light sources different from the dominant one used to obtain the ground truth. This setup mimics realistic scenarios where a single illuminant is not always present. In Table 7, we report results after masking out regions in the testing set that are illuminated by different light sources than the ground truth. Table 8 shows comparisons with other methods on the validation set without masking out regions lit by different illuminations than the dominant light color in the scene.
@@
Qualitative results: Figure 7 presents qualitative results from our method alongside other illuminant estimation methods, namely TLCC [66], FFCC [12], C4 [75], and C5 [7]. We include randomly selected examples representing the top 25%, second quartile, third quartile, and bottom 5% of our results. For FFCC [12] and C5 [7], we show the best result from each method for every example shown in the figure, as we used multiple models for each method—FFCC includes models with and without capture information, and C5 has differently tuned models.

### v2

@@ 5 Conclusion and limitations
While our method represents a promising solution for mobile camera ISPs, its dependency on contextual metadata limits its optimal accuracy to devices that provide geolocation data, which may not available on most DSLR cameras. Additionally, while the contextual metadata is device-independent (i.e., the differences across devices are expected to be minimal), our method relies on additional capture information (i.e., ISO, shutter speed, noise, and SNR stats) and image colors, making it inherently camera-dependent in design. This dependency prevents our trained model from generalizing to new devices without fine-tuning or re-training. However, this issue can be mitigated through color calibration (e.g., [50]), which can be extended to calibrate capture information as well—by performing a pre-processing mapping from the new camera space (for both color and capture information) to the camera space used during training. Additional discussion on cross-camera generalization is provided in the supplementary material.
@@ Appendix B Additional details and results
B.1 Additional details on contextual information
In the main paper, we presented our method, which relies on the “probability” of an image being captured during one of the key solar events (e.g., sunrise, noon, sunset) along with additional capture metadata and color information represented as histograms.
@@ B.1 Additional details on contextual information
B.2 Additional ablation studies
In the main paper, we presented a set of ablation studies to analyze the impact of different input features on our method. Here, we provide additional ablation studies on the validation set with masks applied, as shown in Table 5. By default, our results in Table 5 use the histogram feature, $\mathbf{H}$, and the time-capture feature, $\mathbf{c}$, with noise stats, $\mathbf{n}$.
@@ B.2 Additional ablation studies
B.3 Analysis on outdoor vs. indoor scenes
For outdoor scenes, time-of-day information provides strong cues about the likely range of illuminants. However, it is not as informative for indoor scenes, which are typically illuminated by artificial lights. Therefore, for indoor scenes, image colors and additional capture information are necessary for accurate illuminant estimation.
@@
B.4 Additional quantitative results
In the main paper, we reported results on our testing set without masking out regions illuminated by light sources different from the dominant one used to obtain the ground truth. This setup mimics realistic scenarios where a single illuminant is not always present. In Table 7, we report results after masking out regions in the testing set that are illuminated by different light sources than the ground truth. Table 8 shows comparisons with other methods on the validation set without masking out regions lit by different illuminations than the dominant light color in the scene.
@@ B.4 Additional quantitative results
B.5 Qualitative results
Figure 7 presents qualitative results from our method alongside other illuminant estimation methods, namely TLCC [66], FFCC [12], C4 [75], and C5 [7]. We include randomly selected examples representing the top 25%, second quartile, third quartile, and bottom 5% of our results. For FFCC [12] and C5 [7], we show the best result from each method for every example shown in the figure, as we used multiple models for each method—FFCC includes models with and without capture information, and C5 has differently tuned models.
@@ B.5 Qualitative results
B.6 Cross-camera generalization
In the main paper, we explained that our method is inherently camera-specific by design. However, this limitation can be mitigated through calibration. Here, we present an experimental evaluation of a calibration-based solution to address the camera-specific nature of our approach. Specifically, we calibrate a polynomial mapping for metadata (ISO, shutter speed) and a $3\!\times\!3$ color mapping matrix between our primary camera (Samsung S24 Ultra main camera) and the Samsung S25 Ultra telephoto camera. We evaluate two strategies on 257 test scenes captured by the S25 Ultra telephoto camera:
  1. 1.
    Mapping the S24 Ultra main camera’s training data to the S25 Ultra telephoto camera’s space offline, followed by training on the mapped data.
  2. 2.
    Mapping S25 Ultra telephoto images and metadata online to the S24 Ultra main camera’s space, applying the model trained on the S24 main camera, and then mapping the predicted illuminant back to the S25 Ultra telephoto camera’s space.
Results are shown in Table 9, alongside C4 [75] and C5 [7], both of which are cross-camera methods. None of the methods, including ours, had access to training examples from the target camera (S25 Ultra telephoto).
Table 9: Results on cross-camera generalization. We report the mean angular error for each method on 257 test scenes captured using the S25 Ultra telephoto camera. None of the listed methods were trained on any data from the test camera. For our method, we present results for the model trained on data from the S24 Ultra main camera under three settings: without calibration, with offline calibration, and with online calibration. The best result is highlighted.
Method C4 [75] C5 [7] Ours
w/o calibration w/ offline calibration w/ online calibration
#params (K) 5,116 172 4.8 4.8 4.8
Mean AE (S25-T) 1.61 1.63 2.06 1.70 1.53